OpenStudy (mysticfury):

if these were my set of points, how would i find the slope intercept form? x y 2 –10 4 –16 6 –22 @563blackghost

OpenStudy (ltrout):

You would use Y2-Y1/X2-X1.

OpenStudy (ltrout):

Or you could just count the difference in values on the y side and x side and its Y/X. So you have a 6 difference on the y and 2 on the X so 6/2 = 3/1 = 3 for your slope.

OpenStudy (mysticfury):

okay but then how would i find the y intercept? i know someone said , the y intercept isnt what it says unless the x intercept is 0

OpenStudy (ltrout):

The Y intercept is where the line goes across the Y Axis. If you take the slope of 3 and you subtract that from the -10 you would get (1, -7) and then do it again to get (0, -4) which would be your Y intercept. Meaning the equation for this set of data would be y = 3x-4.

OpenStudy (mysticfury):

oh okay